OpenStudy (anonymous):

x^2+y^2= x - 2y By completing the square find the center and radius of the circle. how would I go about doing this problem?

OpenStudy (anonymous):

\[\left( x^2-x+... \right)+\left( y^2+2y+... \right)=0\] add both sides \[\left( \frac{ coefficient~of~x }{ 2 } \right)^2~\to~complete~the~\square.\] similarly for y

OpenStudy (anonymous):

Center: (1/2,−1) Radius: 5√/ 2 y=−1+−x2+x+1√,−1x2+x+1